What is the path difference of two waves having phase difference 60 degrees?
1
Expert's answer
2012-09-13T11:37:20-0400
Path difference d = ( phase difference * λ) / (2 * pi)
d = ( λ * pi / 3) / (2
* pi) = λ / 6, if these waves have a wavelenth of λ.